OpenStudy (anonymous):

I dont get how to solve this.. Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive

OpenStudy (anonymous):

x = number of votes Maya received: this is what you want to solve for y = total number of votes cast y is given to you = 125,000 Solve for x by plugging in y: x - 62,500 = 0.25(125,000) x - 62,500 = 31,250 x = 93,750 = Number of votes Maya received